### v1

# Detector Response to Gravitational Wave Polarizations in Gravitational Quantum Field Theory

Abstract
We present an analysis of gravitational wave polarization modes within Gravitational Quantum Field Theory (GQFT), a unified theoretical framework reconciling general relativity and quantum field theory. Our study focuses on five fundamental polarization states predicted in GQFT: two tensor (+,×+$\times$+ , ×), two vector (x, y), and one scalar (breathing) mode, focusing on their distinctive detection signatures in space-based interferometers like LISA and Taiji. Using first-order orbital dynamics in the Solar System Barycenter frame, we identify three novel observational features: (1) characteristic interference patterns between polarization modes, (2) distinctive null-point signatures enabling mode discrimination, and (3) sky-position-dependent optimal detection windows. Our approach provides complete sky coverage through polarization mapping while remaining fully compatible with existing mission designs, notably avoiding the need for challenging direct breathing-mode measurements. The results are presented through comprehensive sky maps, offering both theoretical insights into gravitational wave polarization and practical tools for future detector networks. This work establishes a new paradigm for testing fundamental gravity theories through their unique polarization fingerprints, with particular relevance for upcoming multi-messenger gravitational wave astronomy.
1 Introduction
Gravitational waves (GWs) have been extensively observed through various experiments, with polarization playing a crucial role in exploring extended theories of gravity. In the framework of metric-compatible theories, there are up to six possible polarization modes Eardley et al. (1973): two tensor modes (+ and $\times$), two vector modes (x and y), and two scalar modes (breathing and longitudinal). Different gravitational theories predict distinct polarization states. For instance, in general relativity, only the plus and cross modes exist, which have been detected by the LIGO, Virgo, and KAGRA collaborations Abbott et al. (2023). In Brans-Dicke gravity Brans & Dicke (1961); Brans (1962), an additional breathing mode arises due to the presence of a scalar field. Among various alternative theories, this paper focuses on Gravitational Quantum Field Theory (GQFT) Wu (2015, 2016, 2023, 2024).
GQFT was established to reconcile general relativity (GR) and quantum field theory (QFT). It is based on the fundamental principle that the laws of nature are determined by the intrinsic properties of matter’s basic constituents. This principle distinguishes between two types of symmetries: intrinsic symmetries, which are defined by the quantum numbers of quantum fields as elementary particles, and external symmetries, which describe their motion in flat Minkowski spacetimeWu (2015, 2016). According to this framework, the intrinsic spin symmetry SP(1,3) must be localized into a spin gauge symmetry following the gauge symmetry principle.
[…]
The gravigauge field $\chi_{\mu}^{\;a}$ behaves as a Goldstone-boson- type entity, corresponding to a massless graviton. Similar to how gravitational interactions occur for Dirac fermions, the coupling between the gravigauge field $\chi_{\mu}^{\;a}(x)$ and the spin gauge field ${\mathcal{A}}_{\mu}^{ab}$ enables the construction of a spin- gauge-invariant action in GQFT. This structure cannot be replicated using the metric field alone. Importantly, the general linear group symmetry $\mathbb{R}$), which is fundamental to GR, appears as an implicit symmetry within GQFT. A key distinction from GR is that GQFT maintains a flat Minkowski spacetime as its underlying spacetime framework. Interactions between the gravigauge field and the spin gauge/ specified spinor fields introduce non-geometric effects, which result in violations of the equivalence principle observed in GR and give rise to novel physical phenomena Gao et al. (2024). Classically, GQFT recovers GR when the internal spin-gauge sector is deactivated and the new effects stemming from the equivalence principle violation are disregarded.
Beyond its foundational implications, GQFT provides a theoretical framework for exploring a hyperunified field theory capable of unifying all fundamental interactions Wu (2017, 2018a, 2018b, 2021a, 2021b, 2022). Recently, a general theory of the standard model Wu (2025) has been developed within the GQFT framework. This theory integrates the two standard models of particle physics and cosmology, offering new insights into the mysteries of the universe’s dark sector. Specifically, it provides fresh perspectives on understanding the nature of dark matter, the dynamics of the early inflationary universe, and the role of dark energy in the current cosmic expansion.
Numerous studies have been conducted within the framework of GQFT, addressing topics such as inflation Wang et al. (2023), dark matter Wang et al. (2022), and particle physics Huang et al. (2023). The Taiji mission, a space-based gravitational wave (GW) observatory, has been analyzed within the parametrized post-Einsteinian framework to evaluate its capability in detecting different polarization modes. Studies indicate that Taiji can measure both dipole and quadrupole GW emissions Collaboration (2021); Gong et al. (2021); Liu et al. (2020). Similarly, the LISA-Tianqin network has been investigated for its sensitivity to the polarization modes of the stochastic GW background Hu et al. (2024). In a general four-dimensional metric theory of gravity, the detection of additional polarization modes beyond the two predicted by GR would suggest the need for an extended theory of gravitation. The specific polarizations observed could help rule out certain theoretical models Nishizawa et al. (2009a). Both Taiji Hu & Wu (2017) and LISA Danzmann & team (1996) employ a triangular three-detector configuration with equal arm lengths. This paper focuses on the polarization modes in GQFT and their corresponding response characteristics in such a three-detector setup.
The interaction of gravitational waves (GWs) with a detector arm is manifested as fluctuations in the arm length between two detectors in a triangular three-detector configuration. The intensity of the detector’s response depends on the angle between the GW propagation direction and the detector arm. This response is minimal when the GW propagation direction is parallel to the arm and reaches its maximum when the GW direction is perpendicular to the arm. Since gravitational wave sources are distributed across the sky, the angle between the GW propagation direction and the detector arm is not constant. This variation arises due to the orbital motion of space-based detectors, such as Taiji Hu & Wu (2017) and LISA Danzmann & team (1996), around the Sun. As these detectors continuously change their positions and orientations, the angle between the GW propagation direction and the detector arm fluctuates, leading to variations in the detector’s sensitivity to GW signals.
Laser interferometers detect gravitational waves (GWs) by measuring the differential length changes between two detector arms. However, maintaining perfectly equal arm lengths over the entire orbital period of a space-based detector is impractical. This challenge necessitates the use of Time Delay Interferometry (TDI) to account for varying arm lengths. Recent advancements in atomic clock technology offer an alternative approach. By precisely measuring the optical path time difference between detectors, these clocks can infer arm-length variations without relying solely on TDI. This method has sparked significant interest in the development of high-precision optical clocks for space-based GW detectionKolkowitz et al. (2016); Tino et al. (2019); He & Zhang (2020); Wang et al. (2025).
In this paper, we concentrate on investigating novel polarization modes predicted by GQFT that extend beyond GR. Our analysis specifically examines the response of individual detector arms to gravitational wave signals, and will not include considerations of dual-arm interferometric measurements to maintain analytical focus. The paper is organized as follows: it first introduces in §2.1 the polarization modes in metric theories of gravity, then derives in §2.2 the specific polarization modes within the GQFT framework. The analysis further presents in §2.3 the detector response characteristics in the Solar System Barycenter frame, along with a method to distinguish between GQFT and GR based on zero-point distribution patterns, and providing the optimal observational timing windows in §3.
2 Methods
2.1 Polarization in metric theory
The action of GW on test masses can be understood through their tidal effects, which manifest as deviations from geodesic motion. This phenomenon is captured by the Jacobi field equation for geodesic deviation, $\dfrac{\mathop{}\!\mathrm{D}^{2}X^{\alpha}}{\mathop{}\!\mathrm{d}\tau^{2}}=T^{\gamma}\nabla_{\gamma}(T^{\delta}\nabla_{\delta}X^{\alpha})=-R^{\alpha}_{\beta\delta\gamma}X^{\delta}T^{\gamma}T^{\beta}.$ In metric theories of gravity, the polarization state is entirely determined by the electric components of the Riemann tensor Dadhich (2000). When examined in the gravitationally independent flat frame (the natural framework for GQFT), this description simplifies to the form, $\dfrac{\mathop{}\!\mathrm{d}^{2}X^{i}}{\mathop{}\!\mathrm{d}\tau^{2}}=-R^{i}_{0j0}X^{j}$, which describes weak, plane, null gravitational waves Eardley et al. (1973). The symmetry properties of the Riemann tensor reveal that $R^{i}_{0j0}$ possesses six independent components, indicating that gravitational waves can exhibit up to six distinct polarization modes.
For a wave propagating along the z-direction, we define $R_{i0j0}(t)$ as the “driving force matrix” $S_{ij}(t)$, a 3-dimensional symmetric matrix. The polarization modes can be further classified using the Newman-Penrose formalism Newman & Penrose (1962). Working with tetrad indices $(l,n,m,\bar{m})$, we consider tensor contractions with the corresponding null vectors, $X_{efgh}:=X_{\mu\nu\rho\sigma}e^{\mu}f^{\nu}g^{\rho}h^{\sigma}$, where $(e,f,g,h)$ range over the null tetrad basis $(l,n,m,\bar{m})$. Taking a null tetrad frame as follows:
$l=\frac{1}{\sqrt{2}}(\partial_{t}+\partial_{z}),\quad n=\frac{1}{\sqrt{2}}(\partial_{t}-\partial_{z}),$
$m=\frac{1}{\sqrt{2}}(\partial_{x}+i\partial_{y}),\quad\bar{m}=\frac{1}{\sqrt{2}}(\partial_{x}-i\partial_{y}).$ (1)
In the null frame, where the field depends solely on retarded time $(t-z)$, the Riemann tensor satisfies the condition, $R_{efgh,p}=0$. Here, $(e,f,g,h)$ span the null tetrad $(l,n,m,\bar{m})$, while $(p,q)$ range over $(l,m,\bar{m})$. Substituting these into the differential Bianchi identity yields, $R_{ef[pq,n]}={\dfrac{1}{3}R_{efpq,n}=0}$. This implies that, up to trivial constants, we must have: $R_{efpq}=R_{pqef}=0$. Consequently, the only non-vanishing components of the Riemann tensor are of the form $R_{pnqn}$.
Under the null condition, the original 12 NP scalars reduce to just 4 independent quantities: $\Psi_{2},\Psi_{3},\Psi_{4},\Phi_{22}$. Two real scalars $\Psi_{2},\Phi_{22}$, and Two complex scalars $\Psi_{3},\Psi_{4}$. These remaining NP null scalars collectively contain 6 independent degrees of freedom, corresponding exactly to the degrees of freedom in the Riemann tensor for weak, plane, null gravitational waves Hyun et al. (2019):
$\Psi_{2}$ $=\dfrac{1}{6}R_{3030},\quad\quad\Phi_{22}=R_{1010}+R_{2020},\quad\Re(\psi_{3})=\dfrac{1}{2}R_{3010},$
$\Im(\psi_{3})$ $=-\dfrac{1}{2}R_{3020},\quad\Re(\psi_{4})=R_{1010}-R_{2020},\quad\Im(\psi_{4})=-2R_{1020}.$ (2)
Where $\Re$ and $\Im$ represent the real and imaginary part respectively. Then, the driving force matrix can be expanded into a linear combination of polarization bases.
$S(t)=\sum_{A}p_{A}(e_{z},t)E_{A}(e_{z}),$ (3)
where $A$ ranges over (+,×$(+,\times,x,y,b,l),$ covering 6 polarization modes.
In the above expansion, the following bases of polarization matrices are adopted:
$E_{+}(e_{z})$ $:=\begin{bmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{bmatrix},\quad E_{\times}(e_{z}):=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{bmatrix},\quad\quad E_{x}(e_{z}):=\begin{bmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{bmatrix},$
$E_{y}(e_{z})$ $:=\begin{bmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{bmatrix},\quad\quad E_{b}(e_{z}):=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&0\end{bmatrix},\quad E_{l}(e_{z}):=\sqrt{2}\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{bmatrix},$ (4)
where the basis is chosen to be normalized to 2. This normalization condition leads to corresponding polarization amplitudes:
$p_{l}(e_{z},t)$ $=3\sqrt{2}\Psi_{2}=\frac{1}{\sqrt{2}}R_{3030},$
$p_{x}(e_{z},t)$ $=2\Re(\Psi_{3})=R_{3010},$
$p_{y}(e_{z},t)$ $=-2\Im(\Psi_{3})=R_{3020},$
$p_{+}(e_{z},t)$ $=\frac{1}{2}\Re(\Psi_{4})=\frac{1}{2}(R_{1010}-R_{2020}),$
$p_{\times}(e_{z},t)$ $=-\frac{1}{2}\Im(\Psi_{4})=R_{1020},$
$p_{b}(e_{z},t)$ $=\frac{1}{2}\Phi_{22}=\frac{1}{2}(R_{1010}+R_{2020}).$ (5)
2.2 Polarization in GQFT
The polarization properties of GQFT can be systematically addressed within its metric formulation. In this framework, the metric is constructed from the gravigauge field $\chi^{a}_{\mu}$ as that, $\chi_{\mu\nu}=\chi^{a}_{\mu}\chi^{b}_{\nu}\eta_{ab}$. In general, the gravigauge field can always be expressed as the form, $\chi_{\mu}^{\;a}\equiv\eta_{\mu}^{\;a}+h_{\mu}^{\;a}/2$. The corresponding metric takes the form, $\chi_{\mu\nu}=\eta_{\mu\nu}+h_{(\mu\nu)}$, with $h_{(\mu\nu)}\equiv(h_{\mu\nu}+h_{\nu\mu})/2+h_{\mu}^{\;a}h_{\nu}^{\;a}\eta_{ab}/4$. Here, $h_{(\mu\nu)}$ is treated using scalar-vector-tensor decomposition.
Dynamical analysis of GQFT identifies five fundamental degrees of freedom in the gravitational sector Gao et al. (2024). Although this implies the theoretical existence of five gravitational wave polarization modes, only three physical polarizations emerge as observationally significant when analyzing their tidal effects on test masses. The geodesic deviation equation, which governs observable gravitational wave effects, exclusively captures tidal forces. Consequently, the remaining two vector-mode polarizations remain observationally inaccessible within this framework, as they encode intrinsic spin properties and represent manifestations of equivalence principle violations rather than measurable tidal deformations.
For a plane GW propagating along the z-direction, it acts as a perturbation of the metric, implies that the derivative in xy-direction is vanishing. The spin-2 sector reproduces the standard transverse-traceless modes of GR as follows:
[$\ ~{}&\hat{h}_{+}&\hat{h}_{\times}&~{}\\ ~{}&\hat{h}_{\times}&-\hat{h}_{+}&~{}\\ ~{}&~{}&~{}&~{}\end{bmatrix},$ (6)
whilst satisfing $\Box\hat{h}_{ij}=0$.
The spin-1 components of the metric perturbation follow the decomposition: $h_{it}=S_{i},h_{ij}=2\partial_{(i}F_{i)}$. While GQFT dynamics establish the relation between $S_{i}$ and $F_{i}$ with $S_{i}=\partial_{t}F_{i}$. Meanwhile, there is the transverse condition $\partial_{i}F^{i}=0$ for the vector field, implying the traceless condition of the corresponding tensor $h_{ij}$. For plane waves propagating along the z-direction, these constraints reduce the degrees of freedom to just $F_{1}$ and $F_{2}$. The resulting spin-1 sector can therefore be expressed as follows:
$\begin{bmatrix}~{}&\partial_{t}F_{1}&\partial_{t}F_{2}&~{}\\ \partial_{t}F_{1}&~{}&~{}&\partial_{z}F_{1}\\ \partial_{t}F_{2}&~{}&~{}&\partial_{z}F_{2}\\ ~{}&\partial_{z}F_{1}&\partial_{z}F_{2}&~{}\end{bmatrix},$ (7)
whilst satisfying $\Box F_{i}=0.$
The spin-0 components of the metric perturbation are given by: $h_{tt}=0,h_{it}=-\partial_{i}B,h_{ij}=\delta_{ij}\left(-2\psi\right)+2\partial_{i}\partial_{j}E$. GQFT introduces two gauge-invariant variables: $\Phi:=\varphi-\dfrac{1}{2}\partial_{t}B$ and $A:=B+2\partial_{t}E$. The resulting spin-0 part along z-direction is given by,
$\begin{bmatrix}-2\varphi&~{}&~{}&-\partial_{z}B\\ ~{}&-2\psi&~{}&~{}\\ ~{}&~{}&-2\psi&~{}\\ -\partial_{z}B&~{}&~{}&-2\psi+2\partial_{z}^{2}E\end{bmatrix}.$ (8)
GQFT dynamics establish the relations: $\psi=-\dfrac{1}{2}\gamma_{W}\Phi$ and $\partial_{t}A=-(\gamma_{W}-2)\Phi$. All scalar components satisfy the wave equation. The full metric perturbation $h$ is obtained by adding up spin-0 (8), spin-1 (7), and spin-2 (6) components.
This enables the calculation of GW polarization amplitudes in GQFT through the described procedure based on Eq.(5)
$p_{+}$ $=-\frac{1}{2}\partial_{t}^{2}h_{+},\quad p_{\times}=-\frac{1}{2}\partial_{t}^{2}h_{\times},$
$p_{x}$ $=-\frac{1}{2}(\partial_{t}^{2}h_{13}-\partial_{t}\partial_{z}h_{01})\equiv 0,\quad p_{y}=-\frac{1}{2}(\partial_{t}^{2}h_{23}-\partial_{t}\partial_{z}h_{02})\equiv 0,$
$p_{b}$ $=\partial_{t}^{2}\psi,\quad p_{l}=\frac{1}{\sqrt{2}}(\partial_{z}^{2}\varphi-\partial_{t}\partial_{z}^{2}B+\partial_{t}^{2}\psi-\partial_{t}^{2}\partial_{z}^{2}E)\equiv 0.$ (9)
Here, the plus (+) and cross ($\times$) polarization modes in GQFT exactly match those in GR. However, the x and y vector modes completely vanish due to the constraints for the spatial components: $h_{13}=\partial_{z}F_{1},h_{23}=\partial_{z}F_{2}$, and the temporal components: $h_{0i}=S_{i}=\partial_{t}F_{i}$ in GQFT, which together eliminate the spin-1 degrees of freedom from contributing to observable polarizations. Similarly, the longitudinal mode vanishes entirely when considering the gauge-invariant variables, $\Phi:=\varphi-\dfrac{1}{2}\partial_{t}B,A:=B+2\partial_{t}E$, the scalar field relation, $\psi=-\dfrac{1}{2}\gamma_{W}\Phi$ and the constraint equation, $\partial_{t}A=-(\gamma_{W}-2)\Phi$, as well as the wave equation in z direction, $\Box\psi=0$. Consequently, three of the five polarization modes predicted in GQFT become observable by measuring the tidal deformations, they are two transverse-traceless tensor modes (+, $\times$) of GR and one additional breathing mode induced by $\psi$.
Notably, the above observable three polarization modes coincide with those in Brans-Dicke theory. Mapping GQFT’s scalar perturbation to Brans-Dicke’s scalar field reveals identical electric components of the Riemann tensor. However, GQFT exhibits crucial theoretical distinctions. It fundamentally contains vector degrees of freedom absent in Brans-Dicke theory, stemming from its core formulation where the gravigauge field (rather than the metric) serves as the fundamental gravitational field. Furthermore, the GWs in GQFT are strictly transverse, lacking any longitudinal component. This arises because the gravigauge field behaves as a gauge-type Goldstone boson, naturally suppressing longitudinal modes through its transformation properties.
A comprehensive testing framework for GQFT requires extending beyond traditional geodesic deviation measurements to include: spin-gravity coupling effects between test particles and GWs, possible spin-dependent modulation of detector responses and novel detection signatures in spin-polarized measurement systems. This expanded approach may uncover observable manifestations of the currently hidden spin-1 degrees of freedom, potentially revealing the spin-polarization correlations in wave detection, frequency-dependent modulation effects and novel torsion-like couplings in the detector response.
2.3 Response on the detectors
As mentioned above, GW propagation can be expressed as a linear combination of different polarization tensors Nishizawa et al. (2009b),
$\mathbf{h}(t)=h_{+}(t)\varepsilon_{+}+h_{\times}(t)\varepsilon_{\times}+h_{x}(t)\varepsilon_{x}+h_{y}(t)\varepsilon_{y}+h_{b}(t)\varepsilon_{b}+h_{l}(t)\varepsilon_{l}$ (10)
Initially, consider the polarization tensor and the GW propagation direction along $\hat{k}$. By adopting a coordinate system defined by the basis $\{\hat{k},\hat{m},\hat{n}\}$, the corresponding polarization basis can be constructed as follows:
$\varepsilon_{+}$ $=\hat{m}\otimes\hat{m}-\hat{n}\otimes\hat{n},\quad\varepsilon_{\times}=\hat{m}\otimes\hat{n}+\hat{n}\otimes\hat{m},$
$\varepsilon_{x}$ $=\hat{m}\otimes\hat{k}+\hat{k}\otimes\hat{m},\quad\varepsilon_{y}=\hat{n}\otimes\hat{k}+\hat{k}\otimes\hat{n},$
$\varepsilon_{b}$ $=\hat{m}\otimes\hat{m}+\hat{n}\otimes\hat{n},\quad\varepsilon_{l}=\sqrt{2}\hat{k}\otimes\hat{k}.$ (11)
However, actual detection occurs in the Solar System Barycenter (SSB) frame, where the corresponding coordinate system is denoted as $\{\hat{x},\hat{y},\hat{z}\}$. In this frame, GWs require additional projection and mixing compared to the source coordinate system.
Mathematically, these two coordinate systems are related by a rotation, parameterized by Euler angles that form a chart on $SO(3)$. However, their physical interpretations differ: $\{\theta,\varphi\}$ corresponds to the sky location, specifying the GW propagation direction, while $\{\psi\}$ represents the polarization angle, describing the polarization type and its mixing. Thus, in the SSB frame $\{\hat{x},\hat{y},\hat{z}\}$, the GW direction is typically described using the basis $\{\hat{k},\hat{u},\hat{v}\}$,
$\hat{k}$ $=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta),$
$\hat{u}$ $=(\cos\theta\cos\varphi,\cos\theta\sin\varphi,-\sin\theta),$
$\hat{v}$ $=(-\sin\varphi,\cos\varphi,0).$ (12)
This represents an intrinsic rotational difference of $\psi$ from the source coordinate system $\{\hat{k},\hat{m},\hat{n}\}$. The polarization amplitudes are then projected as:
$\hat{m}$ $=\cos\psi\hat{u}+\sin\psi\hat{v},$
$\hat{n}$ $=-\sin\psi\hat{u}+\cos\psi\hat{v}.$ (13)
This projection enables us to consistently treat both the GW polarization response and detector orbital motion within the SSB reference frame.
Assuming the tensor mode is dominated by quadrupole radiation while other modes exhibit dipole-dominated radiation, the waveform incorporating different polarizations is expressed within the parametrized post-Einsteinian framework at 1PN order Epstein & Wagoner (1975); Wagoner & Will (1976); Will (1977); Chatziioannou et al. (2012). For simplicity, we neglect the orbital parameters of the wave source, considering only the waveform amplitudes. Since this work aims to develop a polarization-focused model for distinguishing between GQFT and GR, the analysis remains independent of specific waveform models, thereby maintaining generality.
A comparative analysis of GQFT wave equations reveals that the b-mode and tensor modes differ by a characteristic factor $(1+\gamma_{W})$, where $\gamma_{W}$ represents a fundamental GQFT parameter characterized by the mass ratio $\gamma_{W}=M_{\mathcal{A}}^{2}/\bar{M}_{\kappa}^{2}$ Wu (2015, 2016, 2023); Gao et al. (2024); Wu (2024, 2025). Here, $M_{\mathcal{A}}$ is the mass of spin gauge boson and $\bar{M}_{\kappa}$ denotes the fundamental mass at Planck scale. As previously established, the polarization amplitudes for plus, cross, and other modes are given by: $\dfrac{2\mathcal{M}}{d_{L}}(\mathcal{M}\omega)^{\frac{2}{3}}$, $\dfrac{4\mathcal{M}}{d_{L}}(\mathcal{M}\omega)^{\frac{2}{3}}$ and $\dfrac{\alpha_{A}\mathcal{M}}{d_{L}}(\mathcal{M}\omega)^{\frac{1}{3}}$ respectively Liu et al. (2020); O’Beirne et al. (2019). Here, $d_{L}$ is the luminosity distance, $\mathcal{M}=M\eta^{\frac{3}{5}}$ is the chirp mass with $\eta=\dfrac{m_{1}m_{2}}{M^{2}}$ and $M=m_{1}+m_{2}$ is the total mass.
Turning to detector trajectories, both the LISA and Taiji missions have Keplerian orbital characteristics, maintaining an arm length of 2.5$\sim$ 3 million kilometers. To first order in eccentricity, the spacecraft positions in SSB coordinates can be expressed as Rubbo et al. (2004),
$x(t)$ $=R\cos\alpha+\frac{1}{2}eR\bigl{(}\cos(2\alpha-\beta)-3\cos\beta\bigr{)}$ (14)
$y(t)$ $=R\sin\alpha+\frac{1}{2}eR\bigl{(}\sin(2\alpha-\beta)-3\sin\beta\bigr{)}$ (15)
$z(t)$ $=-\sqrt{3}eR\cos(\alpha-\beta),$ (16)
where $R=1\mathrm{AU}$ represents the radial distance to the guiding center, $e=\dfrac{L}{2\sqrt{3}R}$ denotes the eccentricity, $\alpha=2\pi f_{m}t+\kappa$ is the orbital phase of the guiding center, and $\beta=\dfrac{2\pi n}{3}+\lambda(n=0,1,2)$ describes the relative phase of the spacecraft within the constellation. Setting $\kappa=\lambda=0$ corresponds to the initial ecliptic longitude and orientation of the constellation.
When examining the effect of GWs on the detector, consider a photon emitted from spacecraft $i$ at time $t_{i}$ and received at spacecraft $j$ at time $t_{j}$. The nominal arm length is $\ell_{ij}$, and its variation under low-frequency GWs (where the transfer function approaches unity) can be expressed as Rubbo et al. (2004); Jin & Qiao (2024); Guo et al. (2024)
$\delta\ell_{ij}=\frac{1}{2}\hat{r}_{ij}^{a}(t)\otimes\hat{r}_{ij}^{b}(t)\int^{t_{j}}_{t_{i}}\mathbf{h}_{ab}(t-\hat{k}\cdot\mathbf{x}(t))\mathop{}\!\mathrm{d}t.$ (17)
Here, $a,b$ are abstract indices denoting tensor contractions, and $\hat{r}_{ij}^{a}(t)=\dfrac{\mathbf{x}^{a}_{j}(t_{j})-\mathbf{x}^{a}_{i}(t_{i})}{\ell_{ij}(t_{i})}$ represents the unit vector along the arm direction derived from the orbit (see Eq. (16)).
While the interference response typically measures differential changes between adjacent arms, our analysis primarily focuses on the distinctive breathing mode in GQFT, which is a transverse and isotropic polarization. Since interference subtraction could potentially suppress this mode’s signature, we instead examine single-arm responses through the relative length variation $\dfrac{\delta\ell_{ij}}{\ell_{ij}}$. This approach, combined with a metric perturbation expansion by polarization components, enables isolated examination of each polarization’s contribution.
For clarity and conciseness, we restrict our analysis to periodic detector effects and GWs propagating along the SSB frame’s z-direction. When the GW coordinate system aligns with the detector frame, the arm responses to polarization modes simplify considerably. Let us consider the GW source as a monochromatic signal, with the amplitude response for three polarization modes (generated by compact binary systems like double white dwarfs) given by the following expressions Jin & Qiao (2024),
$A:=\left(\frac{m_{1}m_{2}}{\left(m_{1}+m_{2}\right)^{2}}\right)^{3/5}\left(m_{1}+m_{2}\right)\omega.$ (18)
The detector arm responses to different polarization modes, $h_{ij,+}(t),h_{ij,\times}(t)$ and $h_{ij,b}(t)$, are given by:
$h_{12,+}(t)=\frac{A^{5/3}}{576\omega d_{L}}\Bigg{(}3\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)^{2}\left(\cos^{2}(\theta)\cos^{2}(\varphi)-\sin^{2}(\varphi)\right)+\\ \left(2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+9\right)^{2}\left(\cos^{2}(\theta)\sin^{2}(\varphi)-\cos^{2}(\varphi)\right)+12\sqrt{3}\sin(2\theta)\\ \left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\cos(\varphi)\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)+\frac{3}{2}(\cos(2\theta)+3)\sin(2\varphi)\\ \left(2\sin(4\pi t)+2\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+3\sqrt{3}\right)\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)+12\sin(2\theta)\\ \left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\sin(\varphi)\left(2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+9\right)\\ +144\sin^{2}(\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)^{2}\Bigg{)},$ (19)
$h_{12,\times}(t)=\frac{A^{5/3}}{48\omega d_{L}}\Bigg{(}4\sin(\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\\ \left(\sqrt{3}\left(2\left(\sin(4\pi t-\varphi)+\cos\left(-4\pi t+\varphi+\frac{\pi}{6}\right)\right)+9\sin(\varphi)\right)+9\cos(\varphi)\right)\\ +3\cos(\theta)\left(-12\sin(4\pi t-2\varphi)+\sin(8\pi t-2\varphi)+\sqrt{3}\cos(8\pi t-2\varphi)-9\sin(2\varphi)-9\sqrt{3}\cos(2\varphi)\right)\Bigg{)},$ (20)
and
$h_{12,b}(t)=\frac{A^{4/3}}{1152\omega d_{L}}\Bigg{(}9\sin^{2}(\theta)\cos(2\varphi)\left(\sqrt{3}\sin(8\pi t)+12\cos(4\pi t)-\cos(8\pi t)-9\right)-9\sin^{2}(\theta)\sin(2\varphi)\\ \left(-12\sin(4\pi t)+\sin(8\pi t)+\sqrt{3}\cos(8\pi t)\right)+9\sin^{2}(\theta)\left(16\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)^{2}+9\sqrt{3}\sin(2\varphi)\right)\\ +12\sqrt{3}\sin(2\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\cos(\varphi)\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)\\ +12\sin(2\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\sin(\varphi)\left(2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+9\right)\\ +9(\cos(2\theta)+3)\left(-3\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+\sqrt{3}\sin(4\pi t)+\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)-3\cos(4\pi t)+10\right)\Bigg{)},$ (21)
for the detector Arm12, and
$h_{13,+}(t)=\frac{A^{5/3}}{1152\omega d_{L}}\Bigg{(}9(\cos(2\theta)+3)\cos(2\varphi)\left(\sqrt{3}\sin(8\pi t)-12\cos(4\pi t)+\cos(8\pi t)+9\right)+24\sin(2\theta)\\ \left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\left(2\sqrt{3}\sin\left(4\pi t-\varphi+\frac{\pi}{6}\right)+2\sqrt{3}\cos(4\pi t-\varphi)-9\sin(\varphi)-9\sqrt{3}\cos(\varphi)\right)\\ -9(\cos(2\theta)+3)\sin(2\varphi)\left(12\sin(4\pi t)-\sin(8\pi t)+\sqrt{3}\cos(8\pi t)-9\sqrt{3}\right)\\ +36\sin^{2}(\theta)\left(-4\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+7\sqrt{3}\sin(4\pi t)+11\cos(4\pi t)+2\right)\Bigg{)},$ (22)
$h_{13,\times}(t)=-\frac{A^{5/3}}{48\omega d_{L}}\Bigg{(}4\sin(\theta)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\\ \left(\sqrt{3}(-2\sin(4\pi t-\varphi)-9\sin(\varphi))+2\sqrt{3}\cos\left(4\pi t-\varphi+\frac{\pi}{6}\right)+9\cos(\varphi)\right)\\ +3\cos(\theta)\left(12\sin(4\pi t-2\varphi)-\sin(8\pi t-2\varphi)+\sqrt{3}\cos(8\pi t-2\varphi)+9\sin(2\varphi)-9\sqrt{3}\cos(2\varphi)\right)\Bigg{)},$ (23)
and
$h_{13,b}(t)=\frac{A^{4/3}}{1152\omega d_{L}}\Bigg{(}\left(-2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+9\right)^{2}\left(\cos^{2}(\theta)\sin^{2}(\varphi)+\cos^{2}(\varphi)\right)\\ +3\left(2\sin\left(\pi\left(4t+\frac{1}{6}\right)\right)+2\cos(4\pi t)-9\right)^{2}\left(\cos^{2}(\theta)\cos^{2}(\varphi)+\sin^{2}(\varphi)\right)\\ +3\sin^{2}(\theta)\sin(2\varphi)\left(-2\sin(4\pi t)+2\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)\left(2\sin\left(\pi\left(4t+\frac{1}{6}\right)\right)+2\cos(4\pi t)-9\right)\\ -12\sin(2\theta)\sin(\varphi)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\left(-2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+9\right)\\ +12\sqrt{3}\sin(2\theta)\cos(\varphi)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\left(2\sin\left(\pi\left(4t+\frac{1}{6}\right)\right)+2\cos(4\pi t)-9\right)\\ +144\sin^{2}(\theta)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)^{2}\Bigg{)},$ (24)
for the detector Arm13, and
$h_{23,+}(t)=\frac{A^{5/3}}{48\omega d_{L}}\Bigg{(}-3\sqrt{3}\sin(2\theta)(5\sin(2\pi t)+\sin(6\pi t))\sin(\varphi)+\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)^{2}\\ \left(\cos^{2}(\theta)\sin^{2}(\varphi)-\cos^{2}(\varphi)\right)+3\sin^{2}(4\pi t)\left(\cos^{2}(\theta)\cos^{2}(\varphi)-\sin^{2}(\varphi)\right)\\ -\frac{1}{2}\sqrt{3}(\cos(2\theta)+3)\sin(4\pi t)\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)\sin(2\varphi)\\ +12\sqrt{3}\sin(\theta)\cos(\theta)\sin(2\pi t)\sin(4\pi t)\cos(\varphi)+36\sin^{2}(\theta)\sin^{2}(2\pi t)\Bigg{)}$ (25)
$h_{23,\times}(t)=-\frac{A^{5/3}}{8\omega d_{L}}\Bigg{(}4\sqrt{3}\sin(\theta)\sin(2\pi t)(\cos(4\pi t-\varphi)+3\cos(\varphi))+\cos(\theta)(6\sin(4\pi t-2\varphi)+\sin(8\pi t-2\varphi)-9\sin(2\varphi))\Bigg{)},$ (26)
and
$h_{23,b}(t)=\frac{A^{4/3}}{96\omega d_{L}}\Bigg{(}-3\sqrt{3}\sin(2\theta)(5\sin(2\pi t)+\sin(6\pi t))\sin(\varphi)\\ +\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)^{2}\left(\cos^{2}(\theta)\sin^{2}(\varphi)+\cos^{2}(\varphi)\right)\\ +3\sin^{2}(4\pi t)\left(\cos^{2}(\theta)\cos^{2}(\varphi)+\sin^{2}(\varphi)\right)\\ +\sqrt{3}\sin^{2}(\theta)\sin(4\pi t)\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)\sin(2\varphi)\\ +12\sqrt{3}\sin(\theta)\cos(\theta)\sin(2\pi t)\sin(4\pi t)\cos(\varphi)+36\sin^{2}(\theta)\sin^{2}(2\pi t)\Bigg{)},$ (27)
for the detector Arm23.
In the above formalisms, we neglect polarization mode mixing by setting $\psi=0$. The characteristic amplitude parameter is defined as $A:=\left(\frac{\text{m1}\text{m2}}{(\text{m1}+\text{m2})^{2}}\right)^{3/5}(\text{m1}+\text{m2})\omega$. Here, $m1$ and $m2$ are the component masses of the binary system, and $\omega$ is the GW frequency. Furthermore, $d_{L}$ denotes the luminosity distance, and $\theta$ and $\phi$ represent the angular sky coordinates of the source (polar and azimuthal angles, respectively).
3 Results
We examine the specific case of HM Cancri (RX J0806.3+1527), with the following parameters: sky location: $(\theta,\varphi)=(1.65,2.10)$ radians, component masses: $0.55\mathrm{M}_{\odot}$ and $0.27\mathrm{M}_{\odot}$, GW frequency: $6.22\mathrm{mHz}$, and luminosity distance: $5\mathrm{kpc}$.
The detector response depends on the relative orientation between the detector arms and the GW source, producing both positive and negative values. We therefore consider the absolute value of the relative length change $\delta l/l$ to characterize the signal. This allows us to separately analyze the responses to plus, cross, and breathing polarization modes over a one-year observation period, and to quantify the differences between GR and GQFT predictions.
Figure 1: GW strain response of arm12 to the HM Cancri system at a specific sky location. Left panel: Relative strain components for plus, cross, and breathing polarization modes. Middle panel: Total strain comparison between GR (sum of plus and cross modes) and GQFT (including all three polarizations). Right panel: Temporal evolution over one year (normalized to 0-1), showing GR null points versus GQFT maxima for the source position in the sky map.
Figure 2: Same as Figure 1, showing the gravitational wave response but for arm13.
Figure 3: Gravitational wave response of arm23, following the same presentation as Figure 1.
[…]
This null-point method offers a powerful alternative to direct mode separation, which is observationally challenging. The temporal separation between theory predictions (up to one month in some cases) provides a more robust test than amplitude-based discrimination. Our analysis focuses specifically on distinguishing GR and GQFT through polarization responses, making the wave source details largely irrelevant. However, the annual distribution of null points does encode source information for future study. For far-field GWs, only the polarization response matters.
For every sky location, there exist characteristic time periods when GR predicts null responses while the $b$ mode remains detectable as predicted by GQFT. These optimal observation times could be mapped across the sky, with colors representing the fractional year times of GR nulls and maximal GQFT signals.
4 Conclusion
This study analyzes the polarization modes of Gravitational Quantum Field Theory (GQFT) through a geometrized framework, utilizing the Riemann electric part derived from geodesic deviation. Our calculations demonstrate that three polarization modes: +, $\times$, and breathing out of the five predicted by first-order perturbation theory in GQFT, are observable via geodesic deviation. This discrepancy arises because GQFT incorporates non-geometric elements, with the spin-related gravigauge field serving as the fundamental gravitational field rather than the metric field in General Relativity (GR). Specifically, the spin gauge field plays a crucial role in gravitational interactions. The two vector degrees of freedom are intrinsically linked to the spin structure and only acquire physical significance when considering coupled spin-gravity interactions. This highlights the unique role of spin in mediating gravitational dynamics within the GQFT framework.
We model detector orbits and gravitational wave propagation using first-order approximations within the Solar System Barycenter (SSB) coordinate system. Through the parameterized post-Einsteinian (ppE) formalism, we derive explicit expressions for polarization amplitudes, yielding compact analytical solutions for detector responses in both Taiji and LISA configurations. Using the ultra-compact binary HM Cancri as our benchmark system, we conduct a systematic comparison of polarization responses between GR and GQFT. Our investigation reveals several key findings: Polarization mode interference generates distinctive null points that exhibit marked differences between GR and GQFT predictions. These null points serve as robust discriminants between the two theories, circumventing the observational challenges associated with direct b-mode detection. In the HM Cancri system, we identify consistent temporal separations of approximately two weeks between GR and GQFT null points across both arm12 and arm23 configurations. Additionally, through sky mapping of optimal observation times, we demonstrate that arm23 exhibits a comparatively weaker dependence on source position relative to arm12, highlighting its potential advantages for certain observational scenarios.
The analysis of polarization signatures offers a powerful methodology for investigating gravitational physics beyond standard GR. The GQFT framework naturally incorporates vector modes through its treatment of spin gauge interactions, while alternative theories may introduce distinct signatures via additional spatial dimensions, new fields, or modified spacetime geometry. Future gravitational wave physics and astronomy could benefit significantly from the development of non-traditional detector configurations extending beyond equilateral geometries, the optimization of interferometer arrays specifically tailored for polarization studies, and the integration of multi-messenger astrophysics to cross-correlate polarization signatures with complementary observational data. This methodology provides an alternative approach for testing fundamental theories of gravity by leveraging their characteristic polarization imprints in gravitational wave phenomena.
5 Acknowledgments
This work has been supported in part by the National Key Research and Development Program of China under Grant No.2020YFC2201501, the National Science Foundation of China (NSFC) under Grants No. 12147103 (special fund to the center for quanta-to-cosmos theoretical physics), No. 11821505, the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDB23030100.